OpenStudy (anonymous):

Please help? Medal + Fan! <3 Compare a direct variation to a linear relationship. How are they similar? How are they different?

OpenStudy (solomonzelman):

Okay, so linear equation is any line, like \(\large\color{black}{ y=mx+b }\). Direct variation, is defined by \(\large\color{black}{ y=kx }\) (where "k" is constant) So linear relationship is not necessarily a direct variation -- usually not, (unless b=0).

OpenStudy (solomonzelman):

You can see that in a linear relationship \(\large\color{black}{ y=mx+b }\) where \(\large\color{black}{ b }\) is real. And direct variation, is also, \(\large\color{black}{ y=mx+b }\) but on a condition that \(\large\color{black}{ b=0 }\).

OpenStudy (anonymous):

@KamiBug

OpenStudy (solomonzelman):

basically. Linear relationship is any line Direct variation, is a line that (must be/ is) going through the origin, though (0,0)

OpenStudy (anonymous):

@SolomonZelman Thank you so so much!

OpenStudy (solomonzelman):

Yes, just a couple examples of direct variations. https://www.desmos.com/calculator/addhotulo9 (keep in mind that a direct variation is always a linear relationship to, which is not true the other way. Linear relationship is not necessarily a direct variation.)